OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Part 2: Provide the solution to the system x + 5y = 3 2x + 9y = 4

OpenStudy (anonymous):

x = 3 - 5y 2(3 - 5y) + 9y = 4 6 - 10y +9y = 4 2 = y x = 3 - 5(2) x = -7

OpenStudy (anonymous):

solve the first equation for x via \[x=3-5y\] substitute \[3-5y\] for x in the second equation and write \[2(3-5y)+9y=4\] now you have on variable and one equation. solve for y and then substitute back into either equation to solve for x enough words for you?

OpenStudy (anonymous):

yeah thanks A LOT both of u :]

OpenStudy (anonymous):

jadler has solution so now you have words and solution